OpenStudy (anonymous):

If x2 – 6x – 27 > 0, which statement is correct? A) –9 < x < 3 Reactivate B) –3 < x < 9 C) x < –9 or x > 3 Reactivate D) x < –3 or x > 9

OpenStudy (anonymous):

Again, I believe the answer to this is D but I have a question

OpenStudy (mathstudent55):

The way to solve this is to first solve the corresponding equation, x^2 - 6x - 27 = 0

OpenStudy (anonymous):

I factored the expression and used guess and check to determine the inequality but how would you solve this with out using guess and check to determine the signs for the inequality.

OpenStudy (mathstudent55):

x^2 - 6x - 27 = 0 (x - 9)(x + 3) = 0 x = 9 or x = -3

OpenStudy (mathstudent55):

I'm doing it the way I learned it. First you solve the corresponding equation by simply replacing the inequality sign with an equal sign.

OpenStudy (mathstudent55):

Now that we have x = -3 or x = 9, we draw a number line and plot the two points of interest, -3 and 9. We use open circles bec the inequality sign did not include equal.